The maximum amount of salt that can be dissolved in water at 0C is approximately 350g/L. Calculate the freezing temperature of the saturated solution.
Answer #1

Answer:

The saturated salt solution contains 350 g per Litre of solution.

The density of saturated salt solution = 1.2 g / mL

therefore 1 L of saturated solution weighs 1200 g

it contains 350 g solute

Therefore the solution is a mixture of 350 g solute + 850 g sovent

moles of solute = 350 / 58.45 = 6 moles

mass of solvent = 0.850 Kg

molality of the solution = moles of solute / Kg of solvent

                                 = 6 / 0.850 = 7 mole / Kg

Hence molality = 7

We know that Kf for water = 1.86 K-Kg/mol

Kf, molality, and depression in freezing point are related by the equation

\DeltaT = m Kf

      = 7 mol / Kg x 1.86 K - Kg/Mol

     = 13 K

Hence the freezing temperature of the saturated solution = 0 -13 deg C = -13 deg C

                                                                      or 273 - 13 = 260 K